Hydrated copper (iii) sulphate with mass of 499g is heated. This results in the mass of the anhydrous copper (iii) sulphate being 319g.

Determine the empirical formula for the hydrated copper (iii) sulphate.
1
Expert's answer
2020-06-23T13:01:45-0400

Assume that the formula is "\text{Cu}_2(\text{SO}_4)_3\cdot n\text{H}_2\text{O}".

Find the number of moles of Cu2(SO4)3:


"n_1=\frac{m_1}{\Mu_1}=\frac{319}{2\cdot64+3(32+16\cdot4)}=0.767\text{ mol}."


The amount of substance of water was

"n_2=\frac{m_2}{\Mu_2}=\frac{499-319}{2\cdot1+16}=10\text{ mol}."

From the formula "\text{Cu}_2(\text{SO}_4)_3\cdot n\text{H}_2\text{O}" and the previous calculations, we see the ratio:


"\frac{1}{n}=\frac{n_1}{n_2}=\frac{0.767}{10},\\\space\\
n=\frac{10}{0.767}=13."

Hence, the formula is


"\text{Cu}_2(\text{SO}_4)_3\cdot 13\text{H}_2\text{O}"